OpenStudy (sedatefrog712):

Last question of the night will probably be a really quick question just need some reassurance. Two students, Ann and Max, factored the trinomial 4x^2 − 6x − 4. Ann factored it as 2(x − 2)(2x + 1) and Max factored it as (x − 2)(4x + 2). Indicate which student factored the trinomial completely and which student did not, and explain why.

OpenStudy (sedatefrog712):

@Vocaloid i believe Ann factored it completely but im not completely sure

OpenStudy (sedatefrog712):

this is what i got so far from working it out 2(2x^2-3x-2) 2(2x^2-4-1-2)

OpenStudy (sedatefrog712):

im not completely sure what to do next

Directrix (directrix):

4x^2 − 6x − 4 = 2(2x^2-3x-2) = 2 ( 2x^2 -4x+1x-2) = you want two numbers that multipy to -4 and add to -4

OpenStudy (sedatefrog712):

so -2*2=-4 and -2+-2 = -4

Directrix (directrix):

2 ( 2x^2 -4x+1x - 2) = 2 [ 2x (x -2) + 1* (x - 2)] = (x -2 is a common factor 2 * (x - 2) (2 x + 1)

OpenStudy (sedatefrog712):

ok so max left out the 2 right

Directrix (directrix):

Ann factored it as 2(x − 2)(2x + 1) and Max factored it as (x − 2)(4x + 2). Ann factored it correctly and completely. Max's factoring is correct but is not complete because as you said, he could have factored out another 2.

OpenStudy (sedatefrog712):

oh ok thank you so very much @Directrix it makes a lot more sense now :D

Directrix (directrix):

You are welcome.